OpenStudy (anonymous):

2x – 9y = 6 4x – 3y = –7 Which of the following would be the best first step to solve the system of equations?

OpenStudy (nincompoop):

"which of the following" implies that there are choices to pick from. WHERE ARE THE DAMNED CHOICES?

OpenStudy (shamil98):

WHAT NIN SAID

OpenStudy (anonymous):

A. Multiply the first equation by 2. B. Multiply the second equation by 2. C. Multiply the first equation by –3. D. Multiply the second equation by –3.

OpenStudy (anonymous):

lol oops.

OpenStudy (shamil98):

D

OpenStudy (anonymous):

thanks

OpenStudy (shamil98):

That way you can cross out the y and solve for x.

OpenStudy (solomonzelman):

Eliminate one of the variables, in any of the equations to substitute into the otther. For example eliminate y in the first equation, and whatever y equals to, substitute that for y in the second equation. Understand?

OpenStudy (solomonzelman):

Didn't see the choices.

OpenStudy (anonymous):

For which system of equations is (5, 3) the solution? A. 3x – 2y = 9 3x + 2y = 14 B. x – y = –2 4x – 3y = 11 C. –2x – y = –13 x + 2y = –11 D. 2x – y = 7 2x + 7y = 31

OpenStudy (shamil98):

wickedd, plug them in -_-

OpenStudy (shamil98):

or use wolframalpha >.>

OpenStudy (anonymous):

i know i dont have all too much time /.\

OpenStudy (anonymous):

whats htat

OpenStudy (shamil98):

It takes the same time for me as it would for you. just try -_-

OpenStudy (anonymous):

lol okay

OpenStudy (shamil98):

www.wolframalpha.com

OpenStudy (shamil98):

put your systems in there and it will tell you the solutions..

OpenStudy (anonymous):

i found out its a

OpenStudy (anonymous):

i did the werk

OpenStudy (shamil98):

okay. good job.